ganeshie8 (ganeshie8):

Any cool tricks to find this sum \(\sum \limits_{n=1}^{\infty} \frac{n^2}{10^n} \)

OpenStudy (kainui):

Yeah, I found it out in the last thread. But I'll leave it for someone else to figure out if you want the challenge. =P

OpenStudy (kainui):

I didn't post it there though, so don't bother looking. >=)

ganeshie8 (ganeshie8):

It doesn't look to simplify (by gauss approach) x = 1^2/10 + 2^2/10^2 + 3^2/10^3 + ... 10x = 1^2 + 2^1/10 + 3^2/10^3 + ... ----------------------------------------------------- 9x = 1 + 3/10 + 5/10^2 + 7/10^3 + .... Stuck i guess

OpenStudy (anonymous):

i have an idea, don't know if it is right

OpenStudy (math&ing001):

I think a Taylor-Lagrange would do it.

OpenStudy (anonymous):

write \[f(x)=\sum\frac{n^2x^n}{10^n}\]

OpenStudy (kainui):

You're on the right track @ganeshie8 just notice that you might be able to do the trick again. =)

OpenStudy (anonymous):

You can write that as \[ \sum_{n=1}^\infty \frac{1}{\ln(10)^2} \frac{\partial^2}{\partial \lambda^2} 10^{-\lambda n} \] Interchange the summation and differentiation, then set lambda = 1 at the very end.

ganeshie8 (ganeshie8):

\[\sum_{n=1}^\infty \frac{1}{\ln(10)^2} \frac{\partial^2}{\partial \lambda^2} 10^{-\lambda n} = \frac{1}{\ln(10)^2} \frac{\partial^2}{\partial \lambda^2} \sum_{n=1}^\infty10^{-\lambda n}\] like this ?

ganeshie8 (ganeshie8):

and thats an infinite covnerging geometric series i guess xD

OpenStudy (anonymous):

Yep

OpenStudy (kainui):

I'll go ahead and show how you could derive the most general version of it I can by another method, but it's slower. This method also allows you to easily see how to bootstrap your way up to higher powers of n, like n^3/10^n. differentiate a couple times \[\sum_{n=1}^\infty x^n = \frac{1}{1-x}\]\[\sum_{n=1}^\infty nx^{n-1} = \frac{1}{(1-x)^2}\]\[\sum_{n=1}^\infty n(n-1)x^{n-2} = \frac{2}{(1-x)^3}\] notice that the last one looks like this \[\sum_{n=1}^\infty n^2x^{n-2}-\sum_{n=1}^\infty nx^{n-2} = \frac{2}{(1-x)^3}\] Well that's convenient since our second one above can be changed by a simple division by x. \[\sum_{n=1}^\infty nx^{n-2} = \frac{1}{x(1-x)^2}\] so we get:\[\sum_{n=1}^\infty n^2x^{n-2}=\frac{1}{x(1-x)^2}+ \frac{2}{(1-x)^3}\] Perhaps there's a very general formula suggested by this?

OpenStudy (anonymous):

That is equivalent to using the parameter lambda n times: \[ \frac{1}{\ln(10)^n} \frac{\partial^n}{\partial \lambda^n} \sum_{p=1}^\infty 10^{-\lambda p} = \frac{1}{10\ln(10)^n} \frac{\partial^n}{\partial \lambda ^n} \frac{1}{1-10^{-\lambda}}\]

OpenStudy (anonymous):

Oops- there should be a (-1)^n in there as well.

OpenStudy (kainui):

But this is really the hard way, if you keep going with your first attempt @ganeshie8 you might find it. Just notice that all your numbers are really now turned into odd coefficients so if you move them down by 10 you will have all 2's. So wait did I just derive that basically @Jemurray3 ? Good to know I guess lol.

ganeshie8 (ganeshie8):

woah woah fantastic ! It looks recursive ? we need to find the n/10^n series first right ?

OpenStudy (kainui):

Yeah, but by simply taking the derivative, we had found that as one of our earlier derivatives so it's already there and I plugged it in for us so we didn't have to look at it as a sum anymore. But you're better off continuing: 9x=1+3/10+5/100+7/1000+... 90x=10+3+5/10+7/100+9/1000+... 90x-9x=12+2/10+2/100+2/1000+... Should be straight forward now I think? =P

ganeshie8 (ganeshie8):

brilliant ! got it completely xD

OpenStudy (anonymous):

There is a simpler way to do all these things

ganeshie8 (ganeshie8):

and yes lol looks you have just derived the lagrange-taylory thingy by taking nth derivatives :)

OpenStudy (anonymous):

shall i tell my approach

OpenStudy (kainui):

Go for it, I'm interested haha.

OpenStudy (kainui):

I actually came up with a generalize summation formula that has a variable step size instead of just increasing by 1 each time it can increase by 1/2 or pi or whatever. And the great part is if you plug in 0 you get an integral. If anyone's interested in seeing that, I can show it. The fun part is, it's all algebraic and there's no limit taking.

ganeshie8 (ganeshie8):

sure @No.name im also very curious to see another way to do this !

OpenStudy (anonymous):

If u see the pattern sum((n^0)/(B^n)) is (1 /((B-1)^1)) sum((n^1)/(B^n)) is (B /((B-1)^2)) sum((n^2)/(B^n)) is (B(B+1))/((B-1)^3)

OpenStudy (anonymous):

Take small numbers and evaluate the pattern i bet its right So for this you need to take pattern in patterns

OpenStudy (kainui):

Wait, I don't really know what you're saying @No.name but it sounds pretty much like what I just said above.

ganeshie8 (ganeshie8):

you have got very good eyes @No.name yes it works perfectly !!

OpenStudy (anonymous):

B= 10 So, for n^2 [10(10+1)]/(10-1)^3 110/729

OpenStudy (anonymous):

No need to take partial derivatives

OpenStudy (anonymous):

I disagree. What is the next term? For n = 3?

ganeshie8 (ganeshie8):

haha that ends the nice pattern :(

OpenStudy (anonymous):

try it with any calculus method u would get the same answer as mine

OpenStudy (anonymous):

What is the continuation of that pattern for n = 3?

OpenStudy (anonymous):

or n^3, rather

OpenStudy (anonymous):

It won't

OpenStudy (anonymous):

What do you mean, it won't?

OpenStudy (anonymous):

sum((n^A)/(B^n)) is vA/((B-1)^(A+1))

ganeshie8 (ganeshie8):

I actually came up with a generalize summation formula that has a variable step size instead of just increasing by 1 each time it can increase by 1/2 or pi or whatever. And the great part is if you plug in 0 you get an integral. If anyone's interested in seeing that, I can show it. The fun part is, it's all algebraic and there's no limit taking. you mean n = real number @Kainui ? you're gona explode my head today lol ;=;

OpenStudy (anonymous):

sum((n^A)/(B^n)) is vA/((B-1)^(A+1))

OpenStudy (kainui):

Here's an interesting and slightly useful identity I discovered. Haven't really used it but eh.\[\log \prod_{n=1}^{\infty}f(n)=\sum_{n=1}^\infty \log[f(n)]\]

OpenStudy (anonymous):

So it dosen't end the nice pattern sum((n^A)/(B^n)) is vA/((B-1)^(A+1))

OpenStudy (anonymous):

That's not true.

OpenStudy (anonymous):

what say @ganeshie8

OpenStudy (anonymous):

Also, notice that you can change the step size by replacing n with h*n, where h is the step size. In that case, the sum becomes \[ \frac{(-1)^n}{10h\ln(10)^n} \frac{\partial^n}{\partial \lambda ^n} \frac{1}{1-10^{-\lambda h}} \]

OpenStudy (kainui):

@ganeshie8 I actually don't know if I could generalize it to step sizes of complex numbers... I might try that out but, at any rate here we go: Let's say s is the step size in our summation. I think the important part is that we have our end point and starting point be evenly divisible by this number, otherwise it gets weird... But interesting, kind of like how if x^2 is x multiplied by itself twice then that means x^(1/2) is x multiplied by itself a half a time... But I digress back to the summation. So it builds up recursively and is telescoping. \[\sum_{0}^{a} [(n+s)-n] = [(0+s)-0] +[(s+s)-s] +[(2s+s)-2s] +...\] and it eventually gets to a by taking each term s higher than the last. Think 1 where s is if that helps since that's what you normally would do. The top then simplifies as a telescoping series with the next term subtracting the next just leaving the ends. \[\sum_0^a[(n+s)-n] =a+s\] but this is obviously just \[\sum_0^as =a+s\] pretty straight forward, s is a constant so we can pull it out. So if we plug in numbers as a sanity check, this would be the same as \[\sum_0^a1 =a+1\] which makes sense, we counted 0 right? Also note that if we want a lower bound that's not 0 we can just subtract off one of these sums from itself. That's gonna be more useful for higher order sums. Notice if we plug in 0 then the s in a sense is like dx. \[\sum_0^adx =a\] ok so this looks like an integral, awesome. \[\int\limits _0^adx=a\] ok let's keep going now that we have this.

OpenStudy (anonymous):

A substitution can be made to make it slightly simpler - let \[ 10^{-\lambda h} = e^{x} \rightarrow x = -\lambda h\ln(10)\] Which yields the formula \[ \frac{1}{10^h} \left. \left(\frac{d^n}{dx^n} \frac{1}{1-e^x}\right) \right|_{x = -h\ln(10)}\]

OpenStudy (kainui):

Straightforward, this is a telescoping series so it leaves us with this: \[\sum_0^a[(n+s)^2-n^2]=(a+s)^2\] now we can simplify it on the inside of the sum. \[\sum_0^a 2ns +s^2=\sum_0^a 2ns +\sum_0^as^2=(a+s)^2\] pull out the constants and solve for the one sum with n in it... \[2s \sum_0^a n +s(a+s)=(a+s)^2\]\[s \sum_0^a n =\frac{1}{2}[-s(a+s)+(a+s)^2]\] plug in 1 for s you get \[ \sum_0^a n =\frac{1}{2}[-(a+1)+(a+1)^2]=\frac{(-1+(a+1))(a+1)}{2}=\frac{a(a+1)}{2}\] which is gauss's formula for adding up the numbers to n. Cool! But now let's plug in s=0 to get: \[dx \sum_0^a n =\frac{a^2}{2}\] pretty cool, we are getting the integral all in one! We could also plug in 1/2 and get the sum of all the numbers between 0 and a counting by .5 lol. Continue this process for higher polynomials, pretty interesting.

OpenStudy (kainui):

Some kind of interesting questions, what if your step wasn't a number anymore, but is a function? What if you're stepping into the complex plane and sweeping out a polygon or something? I have no clue, but it works and is pretty fun to play with. But it's sort of a hassle to derive higher order ones since they will all rely on lower order polynomial sums.

OpenStudy (anonymous):

You can replace ten by any base you choose (we could imagine base A, for example) in which case \[ \frac{1}{A^h} \left. \left( \frac{d^n}{dx^n} \frac{1}{1-e^x}\right) \right|_{x=-h\ln(A)} \] That's about as general as I can possibly imagine being. Note that this works if and only if \(|A^{-h}|\) < 1 , or \(|A^h |> 1\). @Kainui if you define the function \(F(x,n) = \frac{d^n}{dx^n} \frac{1}{1-e^x}\), then the sum becomes \[ G(h) = \frac{F(-h\ln(A),n)}{A^h }\] Which can be viewed as a function of h. Analytic continuation into the complex plane is trivial, and would yield a complex function \(G(z)\).

ganeshie8 (ganeshie8):

you folks really make me see my true level in math lol It would definitely take few days for me to comprehend these stuff I guess ! thank you both @Kainui @Jemurray3 for the cool ideas xD

OpenStudy (kainui):

Haha, you're not too far from us I think; I'm just a baby in the math world. If you have any other ideas or thoughts going on you should ask them I'm curious and now I'm in the mood to do more math! lol

OpenStudy (anonymous):

is the formula incorrect

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

why i am ignored, did i say something incorrect or is anyone not here

OpenStudy (anonymous):

it is the erlang series , it is somewhat unpopular

OpenStudy (kainui):

I think it's just cause OS is really glitchy right now.

OpenStudy (anonymous):

Ok

ganeshie8 (ganeshie8):

yeah OS is very laggy :/ lets start a new thread @No.name :)